Grade 11Mechanics

A thick spherical shell has inner radius a & outer b the material has resistivity r. when a potential difference is applies between the innner & outer surfaces, its resistance is ?

P SREENIVASULU
8 Years agoGrade 11

1 Answer

Khimraj
8 Years ago
 
 
We assume a shell element of thickness dp and of radius p.
then resistance is given by
dR = rdp/4πp2
R = (r/4π)\int_{a}^{b}dp/p2 = (r/4π)(1/a – 1/b).
Hope it clears. If you like anser then please approve it.